# MacRAG: Compress, Slice, and Scale-up for Multi-Scale Adaptive Context RAG

### v1

@@ Abstract
Long-context (LC) Large Language Models (LLMs) combined with Retrieval-Augmented Generation (RAG) hold strong potential for complex multi-hop and large-document tasks. However, existing RAG systems often suffer from imprecise retrieval, incomplete context coverage under constrained context windows, and fragmented information caused by suboptimal context construction. We introduce Multi-scale Adaptive Context RAG (MacRAG), a hierarchical retrieval framework that compresses and partitions documents into coarse-to-fine granularities, then adaptively merges relevant contexts through chunk- and document-level expansions in real time. By starting from the finest-level retrieval and progressively incorporating higher-level and broader context, MacRAG constructs effective query-specific long contexts, optimizing both precision and coverage. Evaluations on the challenging LongBench expansions of HotpotQA, 2WikiMultihopQA, and Musique confirm that MacRAG consistently surpasses baseline RAG pipelines on single- and multi-step generation with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o. Our results establish MacRAG as an efficient, scalable solution for real-world long-context, multi-hop reasoning. Our code is available at https://github.com/Leezekun/MacRAG.
@@ 1 Introduction
Large Language Models (LLMs) have substantially advanced question answering and reasoning, yet they frequently exhibit factual gaps or hallucinations when relying solely on internal parameters. Retrieval-Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020) mitigates these pitfalls by equipping LLMs with external evidence from large document collections. Meanwhile, Long Context (LC) LLMs, such as GPT-4o (OpenAI, 2024), Gemini 1.5 (Team et al., 2024), and Llama 3 (Dubey et al., 2024a), push context windows to hundreds of thousands or even millions of tokens, enabling an entire, particularly long context or document to be processed in principle. However, even these extended windows do not fully overcome the “Lost in the Middle” phenomenon (Liu et al., 2024; Xu et al., 2024b), wherein crucial mid-text information is overlooked, and their performance may degrade once length thresholds are surpassed (Leng et al., 2024; Yu et al., 2024).
Within RAG, these issues manifest as three core trade-offs in document processing and retrieval for answer generation:
  1. (1)
    Context Length vs. Focus. Long contexts can improve recall but risk eclipsing critical details (particularly mid-text), whereas shorter contexts narrow in on precision but risk omitting essential evidence.
  2. (2)
    Chunking & Indexing. Finer-grained chunks support precise retrieval but can fragment continuity, while larger chunks preserve coherence but become redundant or unwieldy.
  3. (3)
    Coverage vs. Computation. Contexts across documents for multi-relational queries strain token budgets, latency limits or expensive pre-processing/updates, forcing a balance between broad coverage and efficiency on inference and system operations
These trade-offs become especially acute in enterprise-grade or domain-specific deployments that demand “almost infinite” context for personalization or massive corpora. Moreover, recent iterative RAG approaches Yue et al. (2024); Wang et al. (2025) and agentic RAG pipelines Asai et al. (2024); Jeong et al. (2024); Chen et al. (2025) repeatedly invoke retrieval and generation, compounding inefficiencies unless the underlying retrieval module is robust.
To systematically address these trade-offs, we propose Multi-scale Adaptive Context RAG (MacRAG), a hierarchical framework that unifies top-down offline indexing with bottom-up query-time retrieval. In an offline phase, MacRAG partitions each document into partially overlapping chunks, compresses them (e.g., via abstractive summarization), and further slices these compressed segments for fine-grained indexing via standard overlap chunking. This creates a hierarchical index from document level to fine-grained slices. At query time, MacRAG retrieves the most relevant slices at the finest level to ensure high precision, then progressively scales up to find more borderline contexts. It merges neighboring chunks and performs document-level expansions to construct effective long contexts in real time while controlling context size, addressing precision, coverage, and efficiency trade-offs simultaneously. This hierarchical multi-scale approach is compatible with both single- and multi-step generation methods (Xu et al., 2024a; Zhao et al., 2024) and integrates readily into iterative or agent-based RAG systems. Unlike prior works that improve either retrieval or ranking alone, MacRAG integrates structure-preserving indexing, bottom-up adaptive retrieval, and bounded context assembly for robust multi-hop reasoning. Through extensive evaluations on diverse long-document collections and multi-hop reasoning tasks, MacRAG demonstrates substantial improvements across various generation methods when used as a core RAG module with Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o.
@@ 2 Related Work
2.1 Multi-step Generation with Long Context in RAG
Despite advances in LC LLMs, their integration with RAG systems continues to face the first trade-off discussed in Section 1. Recent research has explored context compression solutions to address the “Lost in the Middle" phenomenon and diminishing returns from excessive context.
Recent approaches have explored two main techniques to effectively handle long contexts via multi-step generation with compression and filtering. Abstractive summarization methods, such as RECOMP-Abst Xu et al. (2024a), generate condensed text by rewriting retrieved passages. Extractive summarization approaches, including RECOMP-Extr Xu et al. (2024a), LLMLingua Jiang et al. (2023a, b), and EXIT Hwang et al. (2024), preserve original text through selective extraction with verification. LongRAG Zhao et al. (2024) offers a different perspective by exploring multiple strategies for utilizing retrieved chunks, comparing direct chunk usage against whole document incorporation and introducing a chain-of-thought-based filter and extractor for selecting relevant information.
However, these post-retrieval multi-step generation approaches for long context face fundamental limitations. First, they rely exclusively on post-retrieval context compression without improving initial retrieval quality, creating an inherent performance ceiling. Second, generation based on compressed rather than original context risks information loss. Third, they incur significant computational overhead through multi-step processing on potentially extensive long context input, i.e., from corpora consisting of excessively long documents. In contrast, MacRAG focuses on the retrieval phase, efficiently constructing precise and comprehensive context inputs with bounded length for generation while minimizing related information loss.
2.2 Hierarchical Retrieval Approaches for RAG
Hierarchical retrieval approaches offer an alternative to post-retrieval compression by capturing longer or broader contexts during the retrieval phase. These methods organize information in hierarchical structures, aiming to improve both precision and recall.
Sarthi et al. (2024) introduce RAPTOR, a tree-based architecture that recursively summarizes clustered text chunks for semantic matching. However, its exclusive focus on semantic similarity presents two key limitations: (1) it often misses important connected relationships that lack direct semantic similarity, hindering multi-hop reasoning capabilities, and (2) its rigid hierarchical structure can fragment interconnected knowledge. SIRERAG Zhang et al. (2025) extends this approach by incorporating relational connectivity alongside semantic similarity in a dual-tree framework. Although this enables more comprehensive knowledge synthesis, it introduces significant computational overhead and depends heavily on accurate proposition extraction and entity-based clustering. The system may miss subtle connections that fall outside its predefined relational model.
Both approaches face substantial computational and practical limitations. Their clustering-based strategies incur high costs during database updates, which is particularly challenging for dynamic knowledge bases. Furthermore, these methods struggle with information overlap across hierarchical levels and lack robust mechanisms to balance semantic similarity with relational relevance. In contrast, MacRAG is more flexible and efficient by preserving the original content structure during indexing and retrieval.
Recent work, such as GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024), adopts hierarchical retrieval through symbolic graphs or sentence-level indexing, which improves coverage but introduces additional overhead. In contrast, MacRAG achieves similar benefits via dense, bottom-up expansion over pre-indexed multi-scale units without symbolic structure. Another line of work bypasses chunking altogether by retrieving entire passages or documents (Qian et al., 2024). While this simplifies retrieval, it reduces granularity control. MacRAG retains a flexible structure through hierarchical slicing and adaptive composition.
@@ 3 Multi-Scale Adaptive Context RAG
We propose Multi-Scale Adaptive Context RAG (MacRAG), a hierarchical multi-scale adaptive retrieval system composed of three main components in sequence. First, the top-down hierarchical indexing processes documents via compression, chunking, and slicing, building a hierarchical index from coarse-grained documents to fine-grained slices. Second, the bottom-up multi-scale adaptive retrieval begins at the finest granularity and progressively traces back to broader contexts, merging relevant information to construct query-adaptive long contexts in real-time. Third, the response generation leverages these merged contexts to produce the final output.
Our primary contribution lies in the retrieval pipeline (phases 1 and 2), which offers a modular foundation for integrating arbitrary response generation techniques in phase 3. Figure 1 concisely illustrates MacRAG, while Algorithm 1 describes its implementation, with mathematical details elaborated in the following sections. The retrieval process is adaptive in the sense that the amount and granularity of retrieved content vary depending on the query’s ambiguity and the initial slice-level relevance scores.
3.1 Document Hierarchical Indexing
Given a document corpus $\mathcal{D}=\{d_{1},\dots,d_{N}\}$, MacRAG performs offline processing to construct a hierarchy through chunking (§3.1.1), compression (§3.1.2), and slicing (§3.1.3), while building a slice-level vector DB at the finest granularity to enable retrieval. Instead of using raw sentences as the base unit, we use partial summaries to reduce redundancy and enhance semantic focus, which helps improve retrieval precision and reduces noise in the multi-scale expansion.
@@ 3.1 Document Hierarchical Indexing
Query $q$, document corpus $\mathcal{D}$, $k_{1}$ (for slice retrieval), $k_{2}$ (for final merged chunks), the number of hops $h$, up-scaling factor $\alpha$
@@
Phase 2: Bottom-up Multi-Scale Adaptive Retrieval
@@ 3.1.1 Chunking
Each document $d\in\mathcal{D}$ is split into partially overlapping chunks,
$\mathcal{C}_{d}=\{c_{1},c_{2},\dots,c_{m}\},$ (1)
where each chunk $c_{i}$ spans approximately 200 to 500 tokens, with a fixed overlap of 10–20 tokens between consecutive chunks. This controlled overlap preserves semantic continuity across chunk boundaries. Although the chunking is token-based by default, it can alternatively be configured at the character level.
@@ 3.1.2 Compression
Each chunk $c_{i}$ is transformed into a more compressed summary $s_{i}$ and thus we obtain a set of compressed summaries $\mathcal{S}_{d}=\{\,s_{1},\dots,s_{m}\}$, where $s_{i}=\mathrm{Compress}(c_{i})$ and $\mathrm{Compress}(\cdot)$ represents a summarization model. It can be either extractive (selecting salient sentences) or abstractive (using a generative model to generate a summary). By default, we use an LLM to generate the abstractive summarization. This step reduces redundancy while retaining core factual and named-entity information.
@@ 3.1.3 Slicing
To facilitate finer-grained retrieval, each summary $s_{i}$ is further split into overlapping slices: $\mathcal{L}_{i}=\bigl{\{}\,\ell_{i,1},\ell_{i,2},\dots,\ell_{i,j},\dots\bigr{\}}$, where each slice $\ell_{i,j}$ typically spans $50\text{-}200$ tokens with partial overlap as in the standard overlap chunking. Each slice is then embedded via $\mathbf{e}_{i,j}=\mathrm{Embed}(\ell_{i,j})$, where $\mathrm{Embed}(\cdot)$ represents a text encoder (e.g., multilingual-e5-large). The slice-level embeddings with metadata (document ID, chunk ID, offsets) are indexed in the vector DB. MacRAG can optionally index chunk and summary embeddings for multi-view retrieval.
3.2 Multi-Scale Adaptive Retrieval
Given an input query $q$, MacRAG implements a multi-scale adaptive retrieval pipeline over the constructed hierarchy through five sequential steps, ultimately selecting the top-$k_{2}$ merged chunks with contextually relevant neighbor expansions for enhanced retrieval precision and recall.
@@ 3.2.1 Initial Slice-level Retrieval
With slices serving as the finest granular units in our hierarchical system and vector DB, we first compute the query-slice relevance scores for each slice $\ell_{i,j}$ in the vector DB:
$r_{q,\ell_{i,j}}=\mathrm{Rel}\!\bigl{(}q,\,\ell_{i,j}\bigr{)},$ (2)
where $\mathrm{Rel}(\cdot,\cdot)$ can be implemented with relevance model, e.g., cosine similarity and $L_{2}$ distance in dense retrieval, sparse retrieval, hybrid retrieval, and so on. We then retrieve the top-$k_{1}$ most relevant slices. This fine-grained retrieval ensures high precision by identifying slices with strong semantic alignment to query $q$.
@@ 3.2.2 Parent Chunk Mapping
Each retrieved slice corresponds to a parent chunk $c_{i}$. To prevent redundancy when multiple retrieved slices originate from the same chunk, we perform a unique mapping:
$\mathcal{C}_{k_{1}}^{{}^{\prime}}=\mathrm{unique}\bigl{(}{c_{i}\mid\ell_{i,\cdot}\in\mathrm{top}\text{-}k_{1}(r_{q,\ell})}\bigr{)},$ (3)
where we identify and consolidate the parent chunks of all retrieved slices and obtain a set of selected chunks used for the next steps.
@@ 3.2.3 Chunk-level Re-Ranking
For each identified unique relevant chunk from the mapping as $c_{i}\in\mathcal{C}_{k_{1}}^{{}^{\prime}}$ to the query, we compute a chunk-level ranking score for the query:
$r_{q,c_{i}}=\mathrm{ReRank}\bigl{(}q,\,c_{i}\bigr{)},$ (4)
utilizing a cross-encoder (e.g., marco-miniLM) or advanced ranking model that evaluates the complete chunk content, mitigating potential “information fragmentation" inherent in slice-level retrieval. The chunks in $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ are then reordered according to their ranking scores to the query $q$.
@@ 3.2.4 Scaling-Up & Document-level Ranking
In this step, we aim to select content segments, which can be used to construct $k_{2}$ precise long contexts with upper bounds in the later step to ensure both relevance and sufficient context coverage for final response generation with LLM. To achieve this, we incorporate top-ranked chunks and their document neighbors, thereby mitigating information fragmentation and enabling multi-hop reasoning. The process consists of the following steps.
Scaled Top Chunk Selection:  Instead of picking exact top-$k_{2}$ chunks, we consider more borderline chunk candidates from the re-ranking stage, since the computed $r_{q,c_{i}}$ may not be perfect. Thus, MacRAG adopts scaling factor $\alpha$ and selects the top-($k_{2}\times\alpha$) chunks from $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ for up-scaling with $r_{q,c_{i}}$ in Equation 4:
@@
where $\alpha\geq 1$ and $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$. $\alpha=1$ does not scale-up the number of borderline candidates, but moderate higher $\alpha$ (e.g., 3 or 4) can provide higher probabilities to include borderline but crucial chunks to improve coverage while maintaining certain-level of qualities for candidates. To balance among precision, recall, and context lengths, $\alpha\in\{2,3,4\}$ would be promising where $(k_{2}\times\alpha)\leq|\mathcal{C}^{{}^{\prime}}_{k_{1}}|$, and we will also demonstrate that selection $\alpha$ is not sensitive, but robust with handling trade-off problems well in Section 4. This approach can also potentially facilitate bridging multi-hop queries.
Document Selection:  We further track one level up from the $(k_{2}\times\alpha)$ chunks $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$ by mapping $c_{i}$ into their parents to identify unique candidates for selecting top-$k_{2}$ documents. We then compute the rank score $r_{q,d}$ of mapped document with re-ranking mechanism leveraging chunks’ rank score $r_{q,c_{i}}$ from all the chunks $c_{i}$ within the document $d$ and $c_{i}$ in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$:
$r_{q,d}\;=\;\mathrm{ReRank_{Doc}}\Bigl{(}\{c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}\}\Bigr{)}.$ (6)
Potential choices for document ranking functions could be mean, max, sum, or an LLM-based ranker. If chunk’s rank score $r_{q,c_{i}}$ is computed by using a cross encoder, then it can be either positive and negative, then $\max({r_{q,c_{i}}})$ would be the base function for document ranking. Based on the computed $r_{q,d}$, we construct the $\mathcal{D}_{q,k_{2}}$ consisting of top-$k_{2}$ documents. We ensure top-$k_{2}$ distinct docs so coverage is not restricted to a single or few docs. If we consider properly higher $\alpha$ (e.g., 3 or 4 depending on the overlapped ratio of original chunks), then there are fewer risks for unfilled top-ranked documents. This is critical when relevant evidence is scattered across multiple documents.
3.2.5 Neighbor Chunk Propagation & Merging

### v2

@@ Abstract
Long-context large language models (LC LLMs) combined with retrieval-augmented generation (RAG) hold strong potential for complex multi-hop and large-document tasks. However, existing RAG systems often suffer from imprecise retrieval, incomplete context coverage under constrained windows, and fragmented information from suboptimal context construction. We introduce Multi-scale Adaptive Context RAG (MacRAG), a hierarchical RAG framework that compresses and partitions documents into coarse-to-fine granularities, then adaptively merges relevant contexts through real-time chunk- and document-level expansions. By initiating with finest-level retrieval and progressively incorporating broader, higher-level context, MacRAG constructs effective query-specific long contexts, optimizing both precision and coverage. Evaluations on challenging LongBench expansions of HotpotQA, 2WikiMultihopQA, and Musique confirm MacRAG consistently surpasses baseline RAG pipelines in single- and multi-step generation using Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o. Our results establish MacRAG as an efficient, scalable solution for real-world long-context, multi-hop reasoning. Our code is available at https://github.com/Leezekun/MacRAG.
@@ 1 Introduction
Large language models (LLMs) have significantly advanced complex reasoning, but they still suffer from factual gaps or hallucinations when relying only on internal parameters (Zhao et al., 2024). Retrieval-Augmented Generation (RAG) addresses this by grounding LLMs in external evidence (Guu et al., 2020; Lewis et al., 2020). Long-context (LC) LLMs such as GPT-4o (OpenAI, 2024), Gemini 1.5 (Team et al., 2024), and Llama 3 (Dubey et al., 2024a) offer large input windows but remain limited. They often miss crucial mid-context information (Liu et al., 2024; Xu et al., 2024b) and their performance degrades at extreme context lengths (Leng et al., 2024; Yu et al., 2024).
These issues in RAG systems give rise to three key trade-offs: (1) Context Length vs. Focus, where longer contexts improve recall but may obscure important details, while shorter ones enhance focus but risk omission of essential evidence; (2) Chunking and Indexing, where fine-grained chunks boost retrieval precision but harm coherence, while coarse chunks preserve structure but introduce redundancy; and (3) Coverage vs. Computation, where broader context improves reasoning but increases token cost and latency, especially in iterative or agentic pipelines (Yue et al., 2024; Asai et al., 2024).
To address these interconnected trade-offs systematically, we propose Multi-scale Adaptive Context RAG (MacRAG). MacRAG integrates top-down offline indexing with bottom-up query-time adaptive retrieval. Offline, documents are partitioned into overlapping chunks, their content compressed via abstractive summarization, and these summaries are further sliced for fine-grained indexing. At query time, MacRAG retrieves precise slices, then adaptively reconstructs the context by merging these into parent chunks, incorporating neighboring chunks, and performing document-level expansions. This constructs effective, query-specific, and length-bounded long contexts, optimizing the balance between precision, coverage, and efficiency.
MacRAG’s unified approach combines structure-preserving indexing, adaptive multi-scale retrieval, and bounded context assembly, offering a distinct, robust solution for complex multi-hop reasoning. Extensive experiments on challenging LongBench (Bai et al., 2023) datasets show significant gains over strong baselines using Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o. This paper thus introduces a novel multi-scale retrieval architecture, empirically validates its significant benefits for demanding QA tasks, and presents an efficient, modular framework for advanced RAG applications.
@@ 2 Related Work
Strategies for RAG with Long Context (LC) LLMs (OpenAI, 2024; Team et al., 2024; Dubey et al., 2024a) largely follow two directions. Post-retrieval context management aims to condense information after initial retrieval by employing methods such as abstractive summarization like RECOMP-Abst (Xu et al., 2024a) or extractive techniques including LLMLingua (Jiang et al., 2023) and RECOMP-Extr (Xu et al., 2024a). While these manage context size, they are bottlenecked by initial retrieval quality, risk information loss, and can be computationally intensive. MacRAG, by contrast, distinctly improves the retrieval phase itself, proactively constructing a high-quality, bounded context.
Orthogonally, hierarchical retrieval approaches organize information during retrieval and ranking for broader coverage. Systems like GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024) use symbolic graphs or sentence-level indexing, which can improve recall but often increase the overhead. RAPTOR (Sarthi et al., 2024) recursively summarizes clustered text chunks, potentially missing non-semantic relations, while SIRERAG (Zhang et al., 2025) further integrates relational connectivity at higher computational cost. LongRAG (Zhao et al., 2024) utilizes entire parent documents of top-ranked chunks alongside a multi-step generation scheme, a strategy that can be inefficient with very long document contexts.
Based on Zhang et al. (2025)’s comparative analysis (Table 4 therein), which shows that GraphRAG underperforms on multi-hop QA while RAPTOR, HippoRAG, and SIRERAG attain similar F1-scores, we focus our evaluations on RAPTOR and LongRAG. Although RAPTOR’s semantic clustering may fragment knowledge and LongRAG’s full-document usage can be costly, both were selected for their promising trade-offs between strong performance and relatively lower overhead compared to other competitive methods, such as the noted underperformance of GraphRAG or the higher operational costs of SIRERAG.
In contrast, MacRAG preserves original document structure through its offline hierarchical indexing and adaptively merges relevant and related contexts via its multi-scale retrieval and ranking at query time. This approach circumvents costly, repeated clustering and the need for explicit symbolic graphs, offering flexible assembly of multi-hop contexts with minimal overhead, and is designed with extendibility towards graph-based enhancements.
@@ 3 Multi-scale Adaptive Context RAG
Query $q$, document corpus $\mathcal{D}$, $k_{1}$ (for slice retrieval), $k_{2}$ (for final merged chunks), the # of hops $h$, up-scaling factor $\alpha$
@@
Phase 2: Bottom-up Multi-scale Adaptive Retrieval and Ranking
@@
We propose Multi-scale Adaptive Context RAG (MacRAG), a hierarchical multi-scale adaptive retrieval system with three sequential components. First, top-down hierarchical indexing compresses documents, partitions them into chunks, and slices them to build a hierarchical index from coarse-grained documents to fine-grained slices with multi-level indices. Second, bottom-up multi-scale adaptive retrieval starts from the finest granularity to ensure precision, and progressively expands to broader contexts by merging relevant information to construct query-specific long contexts dynamically in real time. Third, the response generation leverages these carefully assembled contexts to produce the answer.
Our main contribution lies in the retrieval, ranking, and adaptive input context construction pipeline (phases 1 and 2), which serves as a modular foundation for integrating any response generation method in phase 3. Figure 1 provides an overview of MacRAG, and Algorithm 1 outlines the methods along with equations in subsequent sections.
3.1 Document Hierarchical Indexing
Given a document corpus $\mathcal{D}=\{d_{1},\dots,d_{N}\}$, MacRAG performs offline processing to construct a hierarchy through chunking (§3.1.1), compression (§3.1.2), and slicing (§3.1.3), while building a slice-level vector DB at the finest granularity to empower precise retrieval. This strategy aims to reduce redundancy and sharpen semantic focus, thereby enhancing retrieval precision and minimizing noise during the subsequent multi-scale expansion phase.
@@ 3.1.1 Chunking
Each document $d\in\mathcal{D}$ is split into partially overlapping chunks, $\mathcal{C}_{d}=\{c_{1},c_{2},\dots,c_{m}\}$, where each chunk $c_{i}$ spans approximately 200 to 500 tokens, with a fixed overlap of 10–200 tokens between consecutive chunks. This controlled overlap preserves semantic continuity across chunk boundaries. Although the chunking is token-based by default, it can alternatively be configured at the character level. We applied the similar range of chunk sizes used by LongRAG  (Zhao et al., 2024), more details in Appendix A.5.
@@ 3.1.2 Compression
Each chunk $c_{i}$ is transformed into a more compressed summary $s_{i}$, yielding a set of summaries $\mathcal{S}_{d}=\{s_{1},\dots,s_{m}\}$, where $s_{i}=\mathrm{Compress}(c_{i})$. The $\mathrm{Compress}(\cdot)$ represents a summarization model. It can be either extractive (selecting salient sentences) or abstractive (using a generative model to generate a summary). By default, we use an LLM to generate the abstractive summarization. This step effectively reduces redundancy while retaining core factual information.
@@ 3.1.3 Slicing
To facilitate finer-grained retrieval, each summary $s_{i}$ is further split into overlapping slices: $\mathcal{L}_{i}=\bigl{\{}\,\ell_{i,1},\ell_{i,2},\dots,\ell_{i,j},\dots\bigr{\}}$, where each slice $\ell_{i,j}$ typically spans $50\text{-}200$ tokens with partial overlap as in the standard overlap chunking, and our parameter details are in Appendix A.5. Each slice is then embedded via $\mathbf{e}_{i,j}=\mathrm{Embed}(\ell_{i,j})$, where $\mathrm{Embed}(\cdot)$ represents a text encoder (e.g., multilingual-e5-large). The slice-level embeddings with metadata (document ID, chunk ID, offsets) are indexed in the vector DB. MacRAG can optionally index chunk and summary embeddings for multi-view retrieval.
3.2 Multi-scale Adaptive Retrieval and Ranking
Given an input query $q$, MacRAG implements a multi-scale adaptive retrieval pipeline over the constructed hierarchy through five sequential steps, ultimately selecting the top-$k_{2}$ merged chunks with contextually relevant neighbor expansions for enhanced both precision and recall.
@@ 3.2.1 Initial Slice-level Retrieval
With slices serving as the finest granular units in our hierarchical system and vector DB, we first compute the query-slice relevance scores for each slice $\ell_{i,j}$ in the vector DB: $r_{q,\ell_{i,j}}=\mathrm{Rel}\!\bigl{(}q,\,\ell_{i,j}\bigr{)}$, where $\mathrm{Rel}(\cdot,\cdot)$ can be implemented with relevance model, e.g., cosine similarity and $L_{2}$ distance in dense retrieval, sparse retrieval, hybrid retrieval, and so on. We then retrieve the top-$k_{1}$ most relevant slices. This fine-grained retrieval ensures high precision by identifying slices with strong semantic alignment to the query $q$.
@@ 3.2.2 Parent Chunk Mapping
Each retrieved slice corresponds to a parent chunk $c_{i}$. To prevent redundancy when multiple retrieved slices originate from the same chunk, we perform a unique mapping: $\mathcal{C}_{k_{1}}^{{}^{\prime}}=\mathrm{unique}\bigl{(}{c_{i}\mid\ell_{i,\cdot}\in\mathrm{top}\text{-}k_{1}(r_{q,\ell})}\bigr{)}$, where we identify and consolidate the parent chunks of all retrieved slices and obtain a set of selected chunks used for the next steps.
@@ 3.2.3 Chunk-level Re-Ranking
For each identified unique relevant chunk from the mapping as $c_{i}\in\mathcal{C}_{k_{1}}^{{}^{\prime}}$ to the query, we compute a chunk-level ranking score for the query: $r_{q,c_{i}}=\mathrm{ReRank}\bigl{(}q,\,c_{i}\bigr{)}$, utilizing a cross-encoder (e.g., marco-miniLM) or advanced ranking model that evaluates the complete chunk content, mitigating potential “information fragmentation" inherent in slice-level retrieval. The chunks in $\mathcal{C}_{k_{1}}^{{}^{\prime}}$ are then reordered according to their ranking scores to the query $q$.
@@ 3.2.4 Scaling-Up & Document-level Ranking
This step aims to select optimal content segments for constructing $k_{2}$ precise, bounded long contexts for the final LLM generation, ensuring both relevance and sufficient coverage. It incorporates top-ranked chunks and their context neighbors in the same document to mitigate information fragmentation and support multi-hop reasoning.
@@
Document Selection.  From the scaled set of $(k_{2}\times\alpha)$ chunks in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, we identify their unique source documents. To select the top-$k_{2}$ documents, we compute a document-level rank score $r_{q,d}$ for each candidate document $d$. This score $r_{q,d}$ leverages the chunk-level scores $r_{q,c_{i}}$ of document’s all constituent chunks $c_{i}$ that are also present in $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$: $r_{q,d}=\mathrm{ReRank_{Doc}}\Bigl{(}\{c_{i}\in d\text{~{}and~{}}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}\}\Bigr{)}$. Document ranking functions can include mean, max, sum of chunk scores, or an LLM-based ranker (with $\max(r_{q,c_{i}})$ being a suitable base if using cross-encoder scores, which can be either positive or negative). Based on $r_{q,d}$, we form $\mathcal{D}_{q,k_{2}}$, a set of top-$k_{2}$ distinct documents. Ensuring document diversity, especially with appropriate $\alpha$ values (e.g., 3 or 4 depending on the overlapped ratio of original chunks), mitigates the risk of concentrating on too few sources and is critical when evidence is scattered across multiple documents.
3.2.5 Neighbor Chunk Propagation & Merge for Input Context Construction